### v2

# Rethinking Reward Model Evaluation: Are We Barking up the Wrong Tree?

Abstract
Reward Models (RMs) are crucial for aligning language models with human preferences. Currently, the evaluation of RMs depends on measuring accuracy against a validation set of manually annotated preference data. Although this method is straightforward and widely adopted, the relationship between RM accuracy and downstream policy performance remains under-explored. In this work, we conduct experiments in a synthetic setting to investigate how differences in RM measured by accuracy translate into gaps in optimized policy performance. Our findings reveal that while there is a weak positive correlation between accuracy and downstream performance, policies optimized towards RMs with similar accuracy can exhibit quite different performance. Moreover, we discover that the way of measuring accuracy significantly impacts its ability to predict the final policy performance. Through the lens of Regressional Goodhart’s effect, we identify the existence of exogenous variables impacting the relationship between RM quality measured by accuracy and policy model capability. This underscores the inadequacy of relying solely on accuracy to reflect their impact on policy optimization.
1 Introduction
Reinforcement Learning from Human Feedback (RLHF) (Ibarz et al., 2018; Ouyang et al., 2022) has emerged as a prominent paradigm for aligning Large Language Models (Yang et al., 2024; Dubey et al., 2024). The reward model (Ng & Russell, 2000; Brown & Niekum, 2019; Palan et al., 2019) plays a crucial role in this process by substituting human preferences for model optimization. However, building an RM that fully captures human preferences is highly challenging (Armstrong & Mindermann, 2019; Skalse & Abate, 2023; Lambert et al., 2023). Therefore, the RM can be an imperfect proxy for ideal preferences and cause downstream performance deterioration when optimized against it, known as reward model overoptimization (Gao et al., 2022). This phenomenon, as a result of Goodhart’s law (Karwowski et al., 2023), presents a critical challenge to the RLHF.
The inherent difficulty of constructing an ideal RM require a careful evaluation process to capture its potential negative impace in policy optimization. To date, common practice for evaluating the RM includes directly assessing the optimized policy (Hendrycks et al., 2021; Li et al., 2023) and computing accuracy on a fixed dataset (Lambert et al., 2024). The former, while it can be seen as a final metric in practice, is limited by the cost of optimization. More importantly, it fails to distinguish whether the causes of undesirable behavior arise from the policy optimization process or the reward learning process. The latter, while straightforward, remains the question of whether such evaluation accurately predicts the performance of the downstream optimized policy.
In this work, we systematically investigate the effectiveness of using accuracy in predicting downstream performance. Essentially, the accuracy metric measures the difference between two reward functions, referred to as RM error. In practice, the accuracy quantifies the error between the learned reward model and the human golden reward function. The RM error can lead to potential performance loss between the policies optimized for the proxy RM and the golden reward function. Specifically, we define this performance gap as Policy Regret. Therefore, the goal of RM evaluation is to estimate policy regret through RM error, as depicted in Figure 1.
To investigate how differences in RM measured by accuracy translate into gaps in optimized policy performance, we design a synthetic experiment framework. Due to the inaccessibility of human reward functions, we employ synthetic RM as the golden reward function in our experiments. To effectively collect golden-proxy RM pairs for analysis, we create $N$ different RMs, designating one RM as the golden model and the others as proxy models each time. Based on the above framework, we decompose the issue concerning RM evaluation into three research questions (RQ):
Figure 1: A cartoon of RM error and policy regret.
RQ1: Does the RM error measured by accuracy correlate with the policy regret? We examine the correlation between the accuracy and the policy regret on the widely adopted benchmark dataset (Lambert et al., 2024). We employ optimization algorithms, i.e., best-of-$n$ sampling and policy gradient-based reinforcement learning, for investigation. Our findings reveal a weak positive correlation between the measured accuracy and the policy regret. However, we observe that policies optimized towards reward models within a similar accuracy range can have quite different regrets.
RQ2: How to better measure RM error for policy regret prediction? While we present a positive correlation between the accuracy and the policy regret, there remains room for further enhancement. This naturally raises the question of how to better quantify RM error to achieve a stronger correlation. We begin by investigate the influence of prompt and response distributions. Specifically, we observe that the prompt difference between the RM test dataset and the downstream test dataset will potentially decrease the correlation between the accuracy and the policy regret. Regarding response distribution, we find that the rank of the responses can be more influential rather than the model from which it is sampled. Furthermore, we propose a simple yet effective strategy to enhance the correlation: increasing the number of responses per prompt. Compared to the typical setting, metrics based on more responses reach a higher correlation with the policy regret. We further validate this approach under constraints of equal sample size and annotation budget.
RQ3: What’s the relationship between RM error and Policy Regret? The translation from the RM error to the policy regret can be seen as the result of the reward model overoptimization. Therefore, to accurately predict the policy regret, the quantified RM error should be able to reflect the potential overoptimization. In this work, we derive the expected relationship between the accuracy and the degree of reward overoptimization under the assumptions of Regressional Goodhart’s effect (Manheim & Garrabrant, 2019) and normal distributions of both reward score and noise. However, our findings indicate that reward models with similar accuracy can behave quite differently in terms of overoptimization, suggesting more exogenous variables may take place. This indicates accuracy alone can be an inadequate metric for predicting the downstream performance.
In summary, our work provides deeper understanding into the relationship between RM error and policy regret. This offers valuable for both reward model training and the development of a more robust RLHF algorithm. Our results also highlight the need for a more rigorous benchmark in reward model evaluation and the development of advanced RM assessment tools.
2 Preliminary
Let $\mathcal{X}$ be the set of prompts, and $\mathcal{Y}_{\mathcal{X}}$ be the set of possible responses. A policy $\pi$ is a language model that generates responses $y$ to the prompt $x$ at probability $\pi(y|x)$. The golden reward function assesses the response and gives a score based on their quality $r^{*}:\mathcal{X}\times\mathcal{Y}_{\mathcal{X}}\rightarrow\mathbb{R}$. In practice, the golden reward function $r$ represents the complicated human preference and is generally inaccessible. Instead, a reward model $r$ learned from preference data serves as the proxy of the golden reward function $r^{*}$. Then policy $\pi$ is optimized to maximize the expectation of the reward given by $r$:
$\pi=\max\limits_{\pi}\mathbb{E}_{x\in\mathcal{X},y\sim\pi(\cdot|x)}[r(x,y)]$ (1)
This results in a sub-optimal policy $\pi$ rather than the ideal policy $\pi^{*}$ that directly maximizes the reward given by the golden reward function $r^{*}$.
The workflow of Reinforcement Learning from Human Feedback (RLHF) can be represented as:
  • $\mathcal{A}_{RM}(D_{RM}^{train})\rightarrow r$: Learn a proxy RM $r$ on the preference dataset $D_{RM}^{train}$ by algorithm $\mathcal{A}_{RM}$.
  • $\mathcal{M}(r,r^{*}|D_{RM}^{test})\rightarrow d_{r}$: Evaluate the RM error between proxy reward model $r$ and golden reward function $r^{*}$ by metric $\mathcal{M}$ on the test dataset $D_{RM}^{test}$.
  • $\mathcal{A}_{RL}(\pi_{0},r|D_{RL}^{train})\rightarrow\pi$: Optimize the initial policy $\pi_{0}$ and get policy $\pi$ that maximizes the expectation of reward given by the proxy RM $r$ by the algorithm $\mathcal{A}_{RL}$.
  • $\mathcal{M}^{\prime}(\pi,\pi^{*}|D_{RL}^{test})\rightarrow d_{\pi}$: Measure the regret between the policy $\pi$ and the policy $\pi^{*}$ that optimized directly towards $r^{*}$ by metric $\mathcal{M}^{\prime}$ on the test dataset $D_{RL}^{test}$.
In this work, we essentially focus on the relationship between the error of reward model $d_{r}$ and the regret of policy $d_{\pi}$, which translates our issue into: How should we measure the error $d_{r}$ so as to reach high correlation with the policy regret $d_{\pi}$? Various factors introduced in the RLHF process will potentially influence this relationship. To conduct experiments, we focus on a few interested factors and keep the rest fixed. Specifically, the following factors are considered: The test dataset $D_{RM}^{test}$ on which the reward model $r$ is evaluated; The metric $\mathcal{M}$ used to evaluate the error between the reward models; The test dataset $D_{RL}^{test}$ on which the final policy $\pi$ is evaluated.
(a) Synthetic pipeline of Reinforcement Learning from Human Feedback.
(b) Proxy-golden RM pairs collection for evaluating the correlation between the RM error and policy regret.
Figure 2: Overall experiment framework.
One challenge to perform investigation is the inaccessibility of the golden reward function representing human preferences. To address this difficulty, we follow the previous practice (Gao et al., 2022; 2024) of utilizing another reward model to act as the golden reward function.
Another challenge is to effectively collect the proxy-golden RM pairs for evaluating the correlation between the RM error and policy regret. We implement this by constructing $N=10$ different reward models $r_{i}$ by randomly flipping $\alpha_{i}\%$ of the pairs in the training dataset [^1] Details about dataset construction can be referred to Appendix 8.1. $D_{RM}^{train}$. Then $N$ policies $\pi_{i}$ will be optimized towards the corresponding reward model $r_{i}$. This pipeline is demonstrated in Figure 2(a). This operation will form $N\times N$ golden-proxy RM pairs. As shown in Figure 2(b) , these pair collections can then be used for computing correlation by letting one reward model as the golden reward function and the remaining reward models as proxy reward models.
Due to the inaccessibility of the global optimal policy and the difficulty of constantly controlling optimization pressure, real regret is hard to practically compute under the RLHF context. So, We define Normalised Drop Ratio (NDR) as an estimation of the policy regret. This metric $\mathcal{M}^{\prime}$ quantifies the performance deterioration due to optimization against a proxy reward rather than the golden reward function. It computes the difference between the rewards obtained by the policy $\pi^{*}$ and $\pi$:
$\mathcal{M}^{\prime}_{NDR}(\pi,\pi^{\prime})=\frac{\mathbb{E}_{y\sim\pi(\cdot|x)}[r^{*}(x,y)]-\mathbb{E}_{y\sim\pi_{0}(\cdot|x)}[r^{*}(x,y)]}{\mathbb{E}_{y\sim\pi^{*}(\cdot|x)}[r^{*}(x,y)]-\mathbb{E}_{y\sim\pi_{0}(\cdot|x)}[r^{*}(x,y)]}$ (2)
where the policies $\pi$ and $\pi^{*}$ are optimized against the golden reward function $r^{*}$ and the proxy RM $r$ with exactly the same hyperparameters.
The data from the RewardBench (Lambert et al., 2024) are used to construct the test datasets $D_{RM}^{test}$ and $D_{RL}^{test}$. All reward models are initialized from the Llama-3-instruct-8B model (AI@Meta, 2024) and finetuned by minimizing the negative log-likelihood loss with regularization term:
$\mathcal{L}_{RM}=-\mathbb{E}_{(x,y_{w},y_{l})\sim D_{RM}^{train}}[\log(\sigma(r(x,y_{w})-r(x,y_{l})))]-\mathbb{E}_{(x,y)\sim D_{RM}^{train}}[r(x,y)^{2}]$ (3)
We adopt best-of-$n$ sampling (BoN) and PPO (Schulman et al., 2017) as the algorithm $A_{RL}$ to optimize the initial policy $\pi_{0}$, which is also the Llama-3-instruct-8B model. For the best-of-$n$ sampling, the reward model $r$ picks the response with the highest reward from $n$ candidates. For the PPO algorithm, we follow (Gao et al., 2022) and set the KL penalty in all experiments to be 00.
3 Does the RM error measured by accuracy correlate with the policy regret?
Current research typically assesses reward model errors by computing accuracy on a fixed test set. While this method is straightforward and widely used, few studies have explored whether the accuracy metric correlates with policy regret. In this section, we examine the relationship between measured accuracy and policy regret empirically.
Finding 1:
RM evaluation accuracy is positively related to policy regret, but even with similar RM accuracy, policies can exhibit different levels of regret.
We first investigate the correlation between the accuracy and the policy regret. Original prompts and responses from RewardBench are used to measure accuracy, while the prompts are used for downstream optimization. We assess three traditional correlation coefficients to evaluate how well accuracy metrics benchmark reward models. Additionally, we include the Mean Reciprocal Rank [^2] MRR is calculated by averaging the reciprocal ranks of policies optimized by the reward model with the highest accuracy. The rank of policy is decided by NDR metric. (Voorhees & Tice, 2000) to examine if accuracy can identify the best reward model among multiple candidates. As presented in Table 1, there is a positive relationship between accuracy and policy regret. However, trends depicted in Figure 3(a) and Figure 3(b) illustrate that policy regret can vary considerably even within similar accuracy ranges. Lastly, we observe that accuracy generally correlates more strongly with regret in BoN than in PPO. This is expected, as BoN is a more localized and stable optimization algorithm, making it more predictable by reward model error.
Table 1: The correlation between the accuracy (Acc.) and the policy regret under BoN and PPO setting measured by typical correlation coefficients and Mean Reciprocal Rank (MRR).
{tblr}
width = colspec = Q[77]Q[83]Q[308]Q[173]Q[198]Q[94], cells = c, cell11 = c=2,r=20.16, cell13 = c=40.773, hline1,5 = -0.08em, hline2 = 3-6, hline3 = -, Error / Regret & Relevance
Kendall $\tau$ corr. Pearson corr. Spearman corr. MRR
Acc. BoN 0.6561 0.7520 0.7533 0.6333
Acc. PPO 0.4654 0.6395 0.6102 0.5167
(a) Trend between accuracy $d_{r}^{acc}$ and the policy regret $d^{NDR}_{\pi}$ under BoN setting.
(b) Trend between the accuracy $d_{r}^{acc}$ and the policy regret $d^{NDR}_{\pi}$ under PPO setting.
Figure 3: Trend between the accuracy and policy regret measured by the normalised drop ratio.
4 How to better measure RM error for policy regret prediction?
In the previous section, we examine the positive correlation between accuracy and policy regret. However, there appears to be room for enhancing this correlation, which leads us to the question: how to better quantify the RM error? In this section, we first investigate the influence of prompt and response distribution. Moreover, we explore a straightforward yet effective strategy, i.e., extending responses per prompt. Finally, we validate it under different constraints.
Finding 2:
The rank of responses affects the correlation between accuracy and regret more than the response sampling models.
In the previous experiment, we used the original responses in RewardBench for validation, which were sampled from different models. This naturally raises the question: Can we achieve a higher correlation by sampling responses specifically from the model used for downstream optimization? To examine this question, we construct test datasets with responses solely from different models. We evaluate the correlation between the policy regret and the computed accuracy on these datasets. As shown in Table 2, this approach does not consistently improve upon the original RewardBench dataset. This result suggests that sampling responses from the model used for optimization may not be necessary to achieve a strong correlation between accuracy and policy regret.
Table 2: The correlation between the policy regret and the accuracy on the dataset containing responses sampled from different models. The Origin represents the result on the original RewardBench dataset. The highest results are highlighted in bold, and the second highest result is underlined. The model used for downstream optimization is suffixed with *.
{tblr}
width = colspec = Q[150]Q[206]Q[169]Q[27]Q[206]Q[171], row1 = c, row2 = c, cell11 = r=2, cell12 = c=20.375, cell15 = c=20.377, cell32 = c, cell33 = c, cell34 = c, cell35 = c, cell36 = c, cell42 = c, cell43 = c, cell44 = c, cell45 = c, cell46 = c, cell52 = c, cell53 = c, cell54 = c, cell55 = c, cell56 = c, cell62 = c, cell63 = c, cell64 = c, cell65 = c, cell66 = c, cell72 = c, cell73 = c, cell74 = c, cell75 = c, cell76 = c, hline1,8 = -0.08em, hline2 = 2-3,5-6, hline3,7 = -, Model & BoN PPO
Spearman corr. MRR Spearman corr. MRR
Mistral-7B 0.680±0.042 0.579±0.079 0.580±0.037 0.573±0.073
Llama3-8B* 0.644±0.038 0.598±0.080 0.648±0.047 0.551±0.061
Qwen2-7B 0.680±0.030 0.529±0.074 0.603±0.039 0.599±0.060
Vicuna-7b 0.703±0.035 0.658±0.075 0.527±0.036 0.517±0.068
Origin 0.753 0.633 0.610 0.517
Although sampling from models used for optimization seems unnecessary, we observe varying correlations on datasets constructed by the responses from different models. This suggests potential factors beyond the sampling model, such as the rank of chosen and rejected samples. To examine the assumption, we prepare multiple responses, sort them by their golden rewards, and divide them into bins. We randomly sample chosen and rejected responses from different bins to construct the test datasets with responses of varying ranks. Different trends emerged for BoN and PPO, as demonstrated in Figure 4(a) and Figure 4(b). For BoN, test datasets with chosen samples from average rank bins and rejected samples from lower rank bins exhibited higher correlations. For PPO, datasets with chosen samples from higher rank bins and rejected samples from average rank bins showed a higher correlation. This distinction likely arises from the differing characteristics of the two optimization algorithms. Overall, we conclude the suitable strategy is to choose one response of average rank as a representative sample, along with another response with an appropriate gap for contrast.
(a) The correlation trend on BoN.
(b) The correlation trend on PPO.
Figure 4: The correlation between policy regret and accuracy on datasets constructed from responses of different ranks assessed the by Spearman coefficient. Each grid on the x and y axes represents a sampled bin, with the values below indicating the associated rank.
Finding 3:
The prompt difference between the RM test dataset and downstream test dataset can weaken the accuracy-regret correlation.
The previous experiments assume that the RM and RL test datasets consisted of the same prompts, which may not be true in practice. To investigate the influence of prompt category, we split the test dataset based on the prompt category determined by the original classification in RewardBench. We evaluate accuracy and regret in each category and demonstrate the result in Table 3. We find that accuracy in each category aligns more closely with regret in the corresponding category under the BoN setting. However, this relationship is not present in the PPO setting. This may be because different types of prompts influence each other during PPO optimization.
Table 3: The correlation between policy regret and accuracy on the test dataset composed of prompts from different categories evaluated by the Spearman coefficient. The highest result in each column is bolded, and the highest in each row is underlined.
{tblr}
width = colspec = Q[65]Q[117]Q[150]Q[150]Q[150]Q[150]Q[150], row1 = c, row2 = c, cell11 = c=2,r=20.182, cell13 = c=50.75, cell31 = r=5c, cell33 = c, cell34 = c, cell35 = c, cell36 = c, cell37 = c, cell43 = c, cell44 = c, cell45 = c, cell46 = c, cell47 = c, cell53 = c, cell54 = c, cell55 = c, cell56 = c, cell57 = c, cell63 = c, cell64 = c, cell65 = c, cell66 = c, cell67 = c, cell73 = c, cell74 = c, cell75 = c, cell76 = c, cell77 = c, cell81 = r=5, hline1,13 = -0.08em, hline2 = 3-7, hline3,8 = -, Regret & Accuracy
Chat ChatHard Code Math Safety
BoN Chat 0.529±0.082 0.682±0.058 0.573±0.050 0.557±0.052 0.657±0.044
ChatHard 0.493±0.089 0.682±0.053 0.583±0.038 0.538±0.066 0.655±0.051
Code 0.504±0.095 0.634±0.059 0.717±0.043 0.456±0.060 0.646±0.053
Math 0.288±0.121 0.343±0.080 0.244±0.048 0.610±0.058 0.282±0.059
Safety 0.515±0.093 0.705±0.057 0.521±0.047 0.497±0.067 0.687±0.049
PPO Chat 0.349±0.105 0.500±0.086 0.441±0.054 0.192±0.083 0.611±0.061
ChatHard 0.314±0.115 0.484±0.083 0.434±0.057 0.202±0.087 0.576±0.073
Code 0.384±0.092 0.450±0.080 0.527±0.051 0.371±0.060 0.459±0.053
Math 0.185±0.091 0.275±0.064 0.238±0.050 0.145±0.057 0.313±0.054
Safety 0.359±0.118 0.521±0.062 0.442±0.061 0.332±0.069 0.533±0.053
To further investigate the influence of prompt semantics, we randomly rewrite some prompts in the origin test dataset, creating new test datasets with prompts different from the downstream dataset. We ask GPT-4o (OpenAI, 2024) to perform two different strategies with prompts detailed in Appendix 8.2. One strategy alters the expression without changing the meaning, and the other generates a new prompt within the same category. As shown in Figures 5(a) and 5(b), we find that the correlation is less affected by prompt paraphrasing under BoN. While the correlation on PPO continuously weakens as we paraphrase a larger ratio of prompts.
(a) The correlation trend on BoN.
(b) The correlation trend on PPO.
Figure 5: The correlation between policy regret and accuracy on datasets with different ratios of paraphrased prompts evaluated by the Spearman coefficient and MRR. Different line styles are used to represent datasets paraphrased by different strategies.
Finding 4:
Increasing the number of responses per prompt generally enhances the correlation between measured RM error and regret.
We first examine the strategy of response expansion per prompt by preparing $5$ responses per prompt and employing various metrics [^3] Further details about the computation of these metrics can be found in Appendix 8.3. These metrics include common correlation measures, as well as metrics transported from Information Retrieval. We also explored other metrics for comparison, e.g., expect calibration error and Bo5. Note that some metrics require absolute scoring of responses, which may be inaccessible in practice. As shown in Table 4, metrics evaluated on the dataset with more responses consistently achieve higher correlations in comparison to the accuracy on the dataset with two responses per prompt. Moreover, the $\xi$ correlation (Chatterjee, 2020) generally demonstrated the best performance. However, the ECE metric, despite being relatively common, does not show a significant correlation with regret.
Table 4: The correlation between different metrics and policy regret. The Accuracy-pair stands for the baseline that only samples two responses per prompt and computes the accuracy. Metrics that require an absolute score for each response are suffixed with *.
{tblr}
width = colspec = Q[240]Q[156]Q[131]Q[25]Q[156]Q[131], row1 = c, row2 = c, cell11 = r=2, cell12 = c=20.287, cell15 = c=20.287, cell32 = c, cell33 = c, cell34 = c, cell35 = c, cell36 = c, cell42 = c, cell43 = c, cell44 = c, cell45 = c, cell46 = c, cell52 = c, cell53 = c, cell54 = c, cell55 = c, cell56 = c, cell62 = c, cell63 = c, cell64 = c, cell65 = c, cell66 = c, cell72 = c, cell73 = c, cell74 = c, cell75 = c, cell76 = c, cell82 = c, cell83 = c, cell84 = c, cell85 = c, cell86 = c, cell92 = c, cell93 = c, cell94 = c, cell95 = c, cell96 = c, cell102 = c, cell103 = c, cell104 = c, cell105 = c, cell106 = c, cell112 = c, cell113 = c, cell114 = c, cell115 = c, cell116 = c, cell122 = c, cell123 = c, cell124 = c, cell125 = c, cell126 = c, hline1,13 = -0.08em, hline2 = 2-3,5-6, hline3,12 = -, Metrics & BoN PPO
Spearman corr. MRR Spearman corr. MRR
Pearson corr.* 0.663±0.012 0.657±0.040 0.685±0.016 0.545±0.020
Spearman corr. 0.664±0.015 0.664±0.046 0.680±0.019 0.552±0.030
Kendall $\tau$ corr. 0.665±0.016 0.667±0.042 0.686±0.019 0.558±0.035
Accuracy 0.656±0.015 0.648±0.036 0.671±0.018 0.554±0.025
Bo5* 0.666±0.020 0.663±0.058 0.670±0.024 0.543±0.032
ECE 0.173±0.016 0.364±0.027 0.063±0.014 0.382±0.022
MRR 0.650±0.018 0.646±0.054 0.675±0.022 0.565±0.042
NDCG 0.655±0.032 0.614±0.071 0.658±0.039 0.562±0.053
$\xi$  corr. 0.677±0.026 0.649±0.051 0.688±0.026 0.583±0.047
Accuracy-pair 0.635±0.039 0.586±0.080 0.642±0.047 0.567±0.082
Although measuring more responses per prompt yields a higher correlation, it can be unfair without the control of the total number of samples (prompt-response pairs) in the test dataset. Thus, we further explore the question that should we increase responses or prompts under a constant sample amount [^4] Note that the number of samples approximately translates to the computational cost of the evaluation.. We fix the sample size in the test dataset while varying the number of responses. As shown in Figure 6(a) and Figure 6(b), it is generally more effective to increase the number of responses rather than the number of prompts. However, such bonus decreases with the growing sample size, possibly due to an upper bound in the correlation between regret and measured accuracy.
(a) The correlation trend on BoN.
(b) The correlation trend on PPO.
Figure 6: The correlation between accuracy and policy regret on test datasets with the same number of samples but varying numbers of responses per prompt assessed by the Spearman coefficient.
The annotation cost can be another important factor since building a test dataset with more than two responses can be tricky. To explore this factor, we further examine the question that should we increase responses or prompts under a constant annotation budget. To conduct the experiment, we consider the noise in the annotation process. Among various probability ranking models (Critchlow et al., 1991), we adopt the most commonly used one, i.e., the Bradley-Terry model (Bradley & Terry, 1952). Given a prompt $x$ and two responses $y_{i}$ and $y_{j}$, the probability that $y_{i}$ is preferred to $y_{j}$ is computed by:
$P(y_{i}\succ y_{j}|x)=\frac{1}{1+e^{(r_{j}-r_{i})/\beta}}$ (4)
where $r_{i}$ and $r_{j}$ is the golden reward score given to the response $y_{i}$ and $y_{j}$ in our case. The $\beta$, which can be seen as an indicator of the annotation quality, is set to be $0.5$. We measure the number of pair-wise comparisons by annotation cost. This setting fits the Bradley-Terry model and is also commonly adopted in real-world practice. The annotation process is stimulated by performing a sorting algorithm based on pair-wise comparison.The final results are presented in Figures 7(a) and 7(b). We find that it’s more beneficial to include more responses under the BoN setting with the same annotation budget. However, different trends are observed under the PPO setting.
(a) The correlation trend on BoN.
(b) The correlation trend on PPO.
Figure 7: The relevance between policy regret and the accuracy on test datasets with the same annotation budget while varying responses per prompt are evaluated by the Spearman coefficient.
5 What’s the relationship between RM error and Policy Regret?
In this section, we investigate the relationship between RM error and policy regret. The translation from RM error to the policy regret can be seen as the result of the reward model overoptimization. Therefore, to accurately predict policy regret, the quantified RM error should be able to reveal the potential overoptimization. We first theoretically explore this issue. Then we empirically examine the relationship and analyze the observed optimization dynamics.
Finding 5:
There could be exogenous variables impacting the relationship between RM quality measured by accuracy and policy model optimization.
The translation from RM error to the policy regret can seem as the result of Goodhart’s Laws (Goodhart, 1984). It says that optimizing a less effective metric rather than the golden metric leads to the failure of the system. Such phenomenon is also commonly termed reward model overoptimization (Gao et al., 2022) under the context of RLHF. Among various kinds of Goodhart’s effect leading to the reward model overoptimization, the Regressional Goodhart’s effect (Manheim & Garrabrant, 2019) is the most fundamental one and cannot be avoided. It occurs when the proxy reward model $r$ is equal to the golden reward function $r^{*}$ added with some independent noise $z$, namely:
$r=r^{*}+z$ (5)
Assuming that only the Regressional Goodhart’s effect takes place and the reward score and noise are normally distributed, i.e. $r^{*}\sim\mathcal{N}(0,\sigma_{r}^{2})$ and $z\sim\mathcal{N}(0,\sigma^{2})$, the relationship [^5] Detailed explanation and derivation are provided in Appendix 8.4. between accuracy $d_{r}^{acc}$ and the degree of overoptimization [^6] Assuming normalized reward scores and similar optimization dynamics towards the proxy and true reward models, which generally holds in our BoN experiments, the NDR metric is approximate $d_{\pi}$ times a coefficient. $d_{\pi}$ can be expressed by the following parametric equations:
[…]
where $\Phi$ denotes the cumulative distribution function of standard normal distribution. The equation is too complex to be solved analytically, so we resort to a numerical solution for plotting the relationship. The predicted trend between accuracy and the degree of overoptimization is illustrated in Figure 8(a). We observed that there should be an approximate linear relationship between them.
(a) Expected trend of accuracy $d_{r}^{acc}$ and the degree of overoptimization $d_{\pi}$ by varying the noise $\sigma$.
(b) The actual trend of the accuracy $d_{r}^{acc}$ and the degree of overoptimization $d_{\pi}$ under BoN setting.
Figure 8: Trend between the accuracy $d_{r}^{acc}$ and the degree of overoptimization $d_{\pi}$.
However, the actual trend plotted in Figure 8(b) presents outliers beyond our expected relationship. These discrepancies may stem from additional Goodhart’s effects beyond the Regressional one. We examine the dynamics of golden reward scores when optimizing towards various proxy reward models in Figures 9(a) and 9(b). The distance $d$ is computed as $\sqrt{D_{KL}(\pi||\pi_{0})}$ (Bai et al., 2022). The KL divergence of best-of-$n$ sampling can be computed by $\log n-\frac{n-1}{n}$ (Hilton, 2023). The rewards are computed by an unbiased estimator from (Nakano et al., 2022). The accuracy of these proxy reward models with respect to the golden reward function roughly falls in the range of $(0.7,0.8)$. Despite the close accuracy, we observe distinct behaviors in overoptimization. In the first graph, there is barely any indication of typical overoptimization where the golden reward first increase and then drop. This indicates the Regressional Goodhart’s effect is dominant in this case. In contrast, the second graph shows a noticeable decline in golden reward scores across nearly all proxy models, suggesting the influence of more complex Goodhart’s effects. The differences between these cases suggest the existence of exogenous variables impacting the optimization dynamics.
(a) RM with $\alpha=0\%$ serve as golden reward model.
(b) RM with $\alpha=15\%$ serve as golden reward model.
Figure 9: Trend of overoptimization under BoN setting. The graph plots different golden reward score dynamics when optimizing by different proxy reward models.
6 Related Works
Reinforcement Learning from Human Feedback (RLHF) (Bai et al., 2022; Ouyang et al., 2022; Stiennon et al., 2022) has been a common strategy for the alignment of Large Language Models (Yang et al., 2024; Dubey et al., 2024), in which the reward model plays a crucial role. However, this methodology faces several challenges (Casper et al., 2023; Lang et al., 2024; Armstrong & Mindermann, 2019; Skalse et al., 2022), such as reward model overoptimization (Gao et al., 2022; Lehman et al., 2019), where optimizing policy towards proxy reward model may lead to the ground truth performance deterioration. This phenomenon, as a special case of reward hacking (Krakovna, 2020), can seen as the result of Goodhart’s law (Goodhart, 1984; Zhuang & Hadfield-Menell, 2021).
The RM as an imperfect proxy of the golden preference necessitates the evaluation of whether the reward model accurately captures the real preference (Michaud et al., 2020; Russell & Santos, 2019). Many works consider the measurements to predict the potential consequence due to the difference of reward functions (Ng et al., 1999; Skalse et al., 2023). (Gleave et al., 2021; Skalse et al., 2024) propose the metric quantify the difference between the reward functions and induce regret bounds for optimal policies. However, some works (Zhuang & Hadfield-Menell, 2021; Fluri et al., 2024) propose negative visions of optimization towards imperfect proxy RMs. These works assume the accessibility of the golden reward function, which can be false under the RLHF setting. The common practice under RLHF is to compute accuracy on a fixed preference dataset (Lambert et al., 2024), which remains the question about theoretical or empirical validation of performance prediction.
7 Conclusion
Our study highlights the limitations of evaluating the reward model solely based on accuracy for predicting downstream policy performance. Although there exists a weak positive correlation between accuracy and policy performance, reward models with similar accuracy can result in varying policy outcomes. Moreover, the way of measuring accuracy will significantly impact the prediction performance. Additionally, we identify the presence of exogenous variables influencing the optimization process. Overall, we should take a more cautious attitude about the RM performance indicated by accuracy and develop more advanced techniques for RM evaluation.