### v3

# Rethinking Reward Model Evaluation: Are We Barking up the Wrong Tree?

Abstract
Reward Models (RMs) are crucial for aligning language models with human preferences. Currently, the evaluation of RMs depends on measuring accuracy against a validation set of manually annotated preference data. Although this method is straightforward and widely adopted, the relationship between RM accuracy and downstream policy performance remains under-explored. In this work, we conduct experiments in a synthetic setting to investigate how differences in RM measured by accuracy translate into gaps in optimized policy performance. Our findings reveal that while there is a weak positive correlation between accuracy and downstream performance, policies optimized towards RMs with similar accuracy can exhibit quite different performance. Moreover, we discover that the way of measuring accuracy significantly impacts its ability to predict the final policy performance. Through the lens of the Regressional Goodhart effect, we recognize that accuracy, when used for measuring RM quality, can fail to fully capture the potential RM overoptimization. This underscores the inadequacy of relying solely on accuracy to reflect their impact on policy optimization.
1 Introduction
Reinforcement Learning from Human Feedback (RLHF) (Ibarz et al., 2018; Ouyang et al., 2022) has emerged as a prominent paradigm for aligning Large Language Models (Yang et al., 2024; Dubey et al., 2024). The reward model (Ng & Russell, 2000; Brown & Niekum, 2019; Palan et al., 2019) plays a crucial role in this process by substituting human preferences for model optimization. However, building an RM that fully captures human preferences is highly challenging (Armstrong & Mindermann, 2019; Skalse & Abate, 2023; Lambert et al., 2023). Therefore, the RM can be an imperfect proxy for ideal preferences and cause downstream performance deterioration when optimized against it, known as reward model overoptimization (Gao et al., 2022). This phenomenon, as a result of Goodhart’s law (Karwowski et al., 2023), presents a critical challenge to the RLHF.
The difficulty of constructing an ideal RM require a reliable evaluation to capture potential negative impacts in policy optimization. To date, common practice for evaluating the RM includes directly assessing the optimized policy (Hendrycks et al., 2021; Li et al., 2023) and computing accuracy on a fixed dataset (Lambert et al., 2024). While the former approach serves as a final metric, it is constrained by the cost of optimization and cannot distinguish whether undesirable behaviors stem from the policy optimization process or the reward learning process. The latter approach remains the question of whether such evaluation accurately predicts the performance of the optimized policy.
In this work, we systematically investigate the effectiveness of using accuracy in predicting downstream performance. Essentially, the accuracy metric measures the difference between two reward functions, referred to as RM error. In practice, the accuracy quantifies the error between the learned reward model and the empirical human rewards. The RM error can lead to potential performance loss between the policies optimized for the proxy RM and the golden reward function. Specifically, we define this performance gap as Policy Regret. Therefore, the goal of RM evaluation is to predict policy regret through RM error, as depicted in Figure 1.
To investigate how differences in RM measured by accuracy translate into gaps in optimized policy performance, we design a synthetic experiment framework. Due to the inaccessibility of human reward functions, we employ synthetic RM as the golden reward function in our experiments. To effectively collect golden-proxy RM pairs for analysis, we create $N$ different RMs, designating one RM as the golden model and the others as proxy models each time. Based on the above framework, we decompose the issue concerning RM evaluation into three research questions (RQ):
Figure 1: Translation from RM error to policy regret.
RQ1: Does the RM error measured by accuracy correlate with the policy regret? We examine the correlation between the accuracy and the policy regret on the widely adopted RewardBench dataset(Lambert et al., 2024). We employ optimization algorithms, i.e., best-of-$n$ sampling and policy gradient-based reinforcement learning, for investigation. Our findings reveal a weak positive correlation between the measured accuracy and the policy regret. However, we observe that policies optimized towards reward models within a similar accuracy range can have quite different regrets.
RQ2: How to better measure RM error for policy regret prediction? While we present a positive correlation between the accuracy and the policy regret, there remains room for further enhancement. This naturally raises the question of how to better quantify RM error to achieve a stronger correlation. We begin by investigating the influence of prompt and response distributions. Specifically, we observe that the prompt difference between the RM test dataset and the downstream test dataset will potentially decrease the correlation between the accuracy and the policy regret. Regarding response distribution, we find that the rank of responses evaluated by the golden RM impacts correlation more significantly than the model from which they are sampled. Furthermore, we propose a simple yet effective strategy to enhance correlation: increasing the number of responses per prompt in the test dataset. We validate this approach under constraints of equal sample size and annotation budget, finding that metrics based on more responses achieve a higher correlation with policy regret.
RQ3: What’s the relationship between RM error and Policy Regret? The translation from RM error to policy regret stems from the overoptimization of the reward model. Specifically, optimizing toward an imperfect proxy RM leads to a deterioration in downstream performance. Therefore, to predict the policy regret, the quantified RM error should be able to reflect the potential overoptimization. In this work, we derive the expected relationship between the accuracy and the degree of reward overoptimization under the assumptions of Regressional Goodhart’s effect (Manheim & Garrabrant, 2019) and normal distributions of both reward score and noise. However, our findings indicate that RMs with similar accuracy can behave quite differently in terms of overoptimization. This indicates accuracy alone can be an inadequate metric for predicting the downstream performance.
In summary, our work provides deeper understanding into the relationship between RM error and policy regret. This offers valuable insights for both reward model training and the development of a more robust RLHF algorithm. Our results also highlight the need for a more rigorous benchmark in reward model evaluation and the development of advanced RM assessment tools.
2 Preliminary
Let $\mathcal{X}$ be the set of prompts, and $\mathcal{Y}_{\mathcal{X}}$ be the set of possible responses. A policy $\pi$ is a language model that generates responses $y$ to the prompt $x$ at probability $\pi(y|x)$. The golden reward function assesses the response and gives a score based on their quality $r^{*}:\mathcal{X}\times\mathcal{Y}_{\mathcal{X}}\rightarrow\mathbb{R}$. In practice, the golden reward function $r$ represents the complicated human preference and is generally inaccessible. Instead, a reward model $r$ learned from preference data serves as the proxy of the golden reward function $r^{*}$. Then policy $\pi$ is optimized to maximize the expectation of the reward given by $r$:
$\pi=\max\limits_{\pi}\mathbb{E}_{x\in\mathcal{X},y\sim\pi(\cdot|x)}[r(x,y)]$ (1)
This results in a sub-optimal policy $\pi$ rather than the ideal policy $\pi^{*}$ that directly maximizes the reward given by the golden reward function $r^{*}$.
The workflow of Reinforcement Learning from Human Feedback (RLHF) can be represented as:
  • $\mathcal{A}_{RM}(D_{RM}^{train})\rightarrow r$: Learn a proxy RM $r$ on the preference dataset $D_{RM}^{train}$ by algorithm $\mathcal{A}_{RM}$.
  • $\mathcal{M}(r,r^{*}|D_{RM}^{test})\rightarrow d_{r}$: Evaluate the RM error between proxy reward model $r$ and golden reward function $r^{*}$ by metric $\mathcal{M}$ on the test dataset $D_{RM}^{test}$.
  • $\mathcal{A}_{RL}(\pi_{0},r|D_{RL}^{train})\rightarrow\pi$: Optimize the initial policy $\pi_{0}$ and get policy $\pi$ that maximizes the expectation of reward given by the proxy RM $r$ by the algorithm $\mathcal{A}_{RL}$.
  • $\mathcal{M}^{\prime}(\pi,\pi^{*}|D_{RL}^{test})\rightarrow d_{\pi}$: Measure the regret between the policy $\pi$ and the policy $\pi^{*}$ that optimized directly towards $r^{*}$ by metric $\mathcal{M}^{\prime}$ on the test dataset $D_{RL}^{test}$.
In this work, we essentially focus on the relationship between the error of reward model $d_{r}$ and the regret of policy $d_{\pi}$, which translates our issue into: How should we measure the error $d_{r}$ so as to reach high correlation with the policy regret $d_{\pi}$? Various factors introduced in the RLHF process can potentially influence this relationship. To systematically investigate their impact, we focus on several key factors while keeping other variables fixed. The test dataset $D_{RM}^{test}$ on which the reward model $r$ is evaluated; The metric $\mathcal{M}$ used to evaluate the error between the reward models; The test dataset $D_{RL}^{test}$ on which the final policy $\pi$ is evaluated.
(a) Synthetic pipeline of Reinforcement Learning from Human Feedback.
(b) Proxy-golden RM pairs collection for evaluating the correlation between the RM error and policy regret.
Figure 2: Overall experiment framework.
One challenge in conducting the investigation is the inaccessibility of the golden reward function representing human preferences. To address this difficulty, we follow the previous practice (Gao et al., 2022; 2024) of utilizing another reward model to act as the golden reward function.
Another challenge is to effectively collect the proxy-golden RM pairs for evaluating the correlation between the RM error and policy regret. We implement this by constructing $N=10$ different reward models $r_{i}$ by randomly flipping $\alpha_{i}$ of the pairs in the training dataset [^1] We constructed the data such that if a training set has $\alpha_{i}$ of its labels flipped and another has $\alpha_{j}$ flipped, then $|\alpha_{i}-\alpha_{j}|\%$ of the labels are inconsistent between the two sets. See Appendix 8.1 for dataset details. $D_{RM}^{train}$. Then $N$ policies $\pi_{i}$ will be optimized towards the corresponding reward model $r_{i}$. This pipeline is demonstrated in Figure 2(a). This operation will form $N\times(N-1)$ golden-proxy RM pairs. As shown in Figure 2(b), these pair collections can then be used for computing correlation by letting one reward model as the golden reward function and the remaining reward models as proxy reward models.
Building on prior works that formally define policy regret in the context of general Markov Decision Processes (MDPs) (Karwowski et al., 2023; Fluri et al., 2024), we adapt the concept within the framework of RLHF. Given that the KL divergence between a policy $\pi$ and the initial policy $\pi_{0}$ is $KL(\pi||\pi_{0})=\lambda$, the regret with respect to a golden reward function $r^{*}$ is defined as:
${Reg}_{r^{*}}=\frac{\max\limits_{KL(\pi^{\prime}||\pi_{0})<\lambda}J_{r^{*}}(\pi^{\prime})-J_{r^{*}}(\pi)}{\max\limits_{KL(\pi^{\prime}||\pi_{0})<\lambda}J_{r^{*}}(\pi^{\prime})-\min\limits_{KL(\pi^{\prime}||\pi_{0})<\lambda}J_{r^{*}}(\pi^{\prime})}$ (2)
Here, the function $J_{r^{*}}(\pi)=\mathbb{E}_{y\sim\pi(\cdot|x)}[r^{*}(x,y)]$ represents the expected reward given the policy $\pi$ and the golden reward function $r^{*}$. This value reflects the ratio of the maximum reward gain a policy can achieve to the actual reward gain obtained, considering the constraint of KL divergence.
However, this definition is hard to compute due to the inaccessibility of the optimal policy and the difficulty of controlling KL divergence $\lambda$. Therefore, we propose Normalised Drop Ratio (NDR) as a computable estimation as follows:
$\mathcal{M}^{\prime}_{NDR}(\pi,\pi^{*})=\frac{J_{r^{*}}(\pi)-J_{r^{*}}(\pi_{0})}{J_{r^{*}}(\pi^{*})-J_{r^{*}}(\pi_{0})}$ (3)
where the policies $\pi$ and $\pi^{*}$ are achieved by optimizing initial policy $\pi_{0}$ against the golden reward function $r^{*}$ and the proxy RM $r$ with exactly the same hyperparameters. This metric quantifies the performance deterioration resulting from optimizing against a proxy RM $r$ instead of the golden reward function $r^{*}$. It also accounts for variations due to different scales among different RMs.
The data from the RewardBench (Lambert et al., 2024) are used to construct the test datasets $D_{RM}^{test}$ and $D_{RL}^{test}$. All RMs are initialized from the Llama-3-instruct-8B model (AI@Meta, 2024) and finetuned by minimizing the negative log-likelihood loss with regularization term (Hou et al., 2024):
$\mathcal{L}_{RM}=-\mathbb{E}_{(x,y_{w},y_{l})\sim D_{RM}^{train}}[\log(\sigma(r(x,y_{w})-r(x,y_{l})))]-\mathbb{E}_{(x,y)\sim D_{RM}^{train}}[r(x,y)^{2}]$ (4)
We adopt best-of-$n$ sampling (BoN) and PPO (Schulman et al., 2017) as the algorithm $A_{RL}$ to optimize the initial policy $\pi_{0}$, which is also the Llama-3-instruct-8B model. For the best-of-$n$ sampling, the reward model $r$ picks the response with the highest reward from $n$ candidates. For the PPO algorithm, we follow (Gao et al., 2022) and set the KL penalty in all experiments to be 00.
3 Does the RM error measured by accuracy correlate with the policy regret?
Current research typically assesses reward model errors by computing accuracy on a fixed test set. Although widely used, it remains unclear whether the accuracy correlates with policy regret. In this section, we evaluate the accuracy and policy regret using the RewardBench (Lambert et al., 2024) dataset. We perform deduplication on the prompts, leaving 2,733 distinct prompts.
Finding 1:
RM evaluation accuracy is positively related to policy regret, but even with similar RM accuracy, policies can exhibit different levels of regret.
We first investigate the correlation between accuracy and policy regret. Original RewardBench prompts and responses are used to measure accuracy, and the prompts are used for downstream optimization. We calculate the correlations between RM error, measured by accuracy, and policy regret, measured by NDR, in Table 1. As presented in Figure 3(a) and Figure 3(b), there is a positive relationship between accuracy and policy regret. However, trends illustrate that policy regret can vary considerably even within similar accuracy ranges. Lastly, we observe that accuracy generally correlates more strongly with regret in BoN than in PPO. This is expected, as BoN is a more localized and stable optimization algorithm, making it more predictable by reward model error.
Table 1: The correlation between the accuracy (Acc.) and the policy regret under BoN and PPO measured by typical correlation coefficients and Mean Reciprocal Rank (MRR). MRR is calculated by averaging the reciprocal ranks of policies optimized by the RM with the highest accuracy.
{tblr}
cells = c, cell11 = c=2,r=2, cell13 = c=4, hline1,5 = -0.08em, hline2 = 3-6, hline3 = -,
Error / Regret & Relevance
Kendall $\tau$ corr. Pearson corr. Spearman corr. MRR
Acc. BoN 0.6561 0.7520 0.7533 0.6333
Acc. PPO 0.4654 0.6395 0.6102 0.5167
(a) Trend between accuracy $d_{r}^{acc}$ and the policy regret $d^{NDR}_{\pi}$ under BoN setting.
(b) Trend between the accuracy $d_{r}^{acc}$ and the policy regret $d^{NDR}_{\pi}$ under PPO setting.
Figure 3: Trend between the accuracy and policy regret measured by the normalised drop ratio.
4 How to better measure RM error for policy regret prediction?
In the previous section, we examine the positive correlation between accuracy and policy regret. However, there appears to be room for enhancing this correlation, which leads us to the question: how to better quantify the RM error? In this section, we first investigate the influence of prompt and response distribution. Moreover, we explore a straightforward yet effective strategy, i.e., extending responses per prompt. Finally, we validate it under different constraints.
Finding 2:
The rank of responses affects the correlation between accuracy and regret more than the response sampling models.
In the previous experiment, we used the original responses in RewardBench for validation, which were sampled from different models. This naturally raises the question: Is it possible to achieve a higher correlation by sampling responses exclusively from the model used for downstream optimization? To examine this question, we construct multiple test datasets, each containing responses exclusively sampled from a single downstream model.. We then evaluate the correlation between policy regret and computed accuracy. As shown in Table 2, this approach does not consistently improve upon the original RewardBench dataset. This result suggests that sampling responses from the model used for optimization may not be necessary to achieve a strong correlation.
Table 2: The correlation between policy regret and accuracy on datasets with responses sampled from different models. The Origin represents the result on the RewardBench dataset. The highest results are highlighted in bold. The model used for downstream optimization is suffixed with *.
Model BoN PPO
Spearman corr. MRR Spearman corr. MRR
Mistral-7B 0.680±0.042 0.579±0.079 0.580±0.037 0.573±0.073
Llama3-8B* 0.644±0.038 0.598±0.080 0.648±0.047 0.551±0.061
Qwen2-7B 0.680±0.030 0.529±0.074 0.603±0.039 0.599±0.060
Vicuna-7b 0.703±0.035 0.658±0.075 0.527±0.036 0.517±0.068
Origin 0.753 0.633 0.610 0.517
While it seems unnecessary to sample from models used for optimization, we observe varying correlations in datasets constructed with responses by different models. This indicates that other factors, such as the rank of chosen and rejected samples, may be at play. To examine this assumption, we prepare multiple responses for all prompts. Each time, we sort the responses by golden rewards given by one RM, and divide them into bins. We then randomly sample accepted and rejected responses from different bins to construct test datasets with responses of varying ranks. The accuracy on these test datasets is used to assess the correlation with policy regrets. As demonstrated in Figure 4(a) and Figure 4(b), different trends emerged for BoN and PPO. For BoN, test datasets with chosen samples from mid-rank bins and rejected samples from lower-rank bins exhibited higher correlations. In contrast, for PPO, datasets with chosen samples from higher-rank bins and rejected samples from mid-rank bins showed a stronger correlation. This distinction likely stems from the characteristics of the two optimization algorithms. With PPO, the policy is optimized more aggressively, causing responses to rank higher under the golden RM, making the consistency between the proxy RM and the golden RM on high-reward samples more critical. Conversely, for BoN, where the degree of policy optimization is less intense, consistency on mid-reward samples can be more instructive.
(a) The correlation trend on BoN.
(b) The correlation trend on PPO.
Figure 4: The correlation between policy regret and accuracy on datasets constructed from responses. Each grid on the x and y axes represents a sampled bin, with the values below indicating the corresponding rank. For example, the $5$ on the y-axis indicates that the chosen responses are sampled from those ranked between $5$ and $10$.
Finding 3:
The prompt difference between the RM test dataset and downstream test dataset can weaken the accuracy-regret correlation.
The previous experiments assume that the RM and RL test datasets consisted of the same prompts, which may not be true in practice. To investigate the influence of prompt category, wWe split the test dataset based on the prompt categories determined by the original classifications in RewardBench. We evaluate accuracy and regret in each category and demonstrate the results in Table 3. We find that accuracies in each category align more closely with regret in the corresponding category under the BoN setting. However, this relationship is not present in the PPO setting. This may be because different types of prompts influence each other during PPO optimization.
Table 3: The correlation between policy regret and accuracy on the test dataset composed of prompts from different categories evaluated by the Spearman coefficient. The highest result in each column is bolded, and the highest in each row is underlined.
{tblr}
width = colspec = Q[65]Q[117]Q[150]Q[150]Q[150]Q[150]Q[150], row1 = c, row2 = c, cell11 = c=2,r=20.182, cell13 = c=50.75, cell31 = r=5c, cell33 = c, cell34 = c, cell35 = c, cell36 = c, cell37 = c, cell43 = c, cell44 = c, cell45 = c, cell46 = c, cell47 = c, cell53 = c, cell54 = c, cell55 = c, cell56 = c, cell57 = c, cell63 = c, cell64 = c, cell65 = c, cell66 = c, cell67 = c, cell73 = c, cell74 = c, cell75 = c, cell76 = c, cell77 = c, cell81 = r=5, hline1,13 = -0.08em, hline2 = 3-7, hline3,8 = -, Regret & Accuracy
Chat ChatHard Code Math Safety
BoN Chat 0.529±0.082 0.682±0.058 0.573±0.050 0.557±0.052 0.657±0.044
ChatHard 0.493±0.089 0.682±0.053 0.583±0.038 0.538±0.066 0.655±0.051
Code 0.504±0.095 0.634±0.059 0.717±0.043 0.456±0.060 0.646±0.053
Math 0.288±0.121 0.343±0.080 0.244±0.048 0.610±0.058 0.282±0.059
Safety 0.515±0.093 0.705±0.057 0.521±0.047 0.497±0.067 0.687±0.049
PPO Chat 0.349±0.105 0.500±0.086 0.441±0.054 0.192±0.083 0.611±0.061
ChatHard 0.314±0.115 0.484±0.083 0.434±0.057 0.202±0.087 0.576±0.073
Code 0.384±0.092 0.450±0.080 0.527±0.051 0.371±0.060 0.459±0.053
Math 0.185±0.091 0.275±0.064 0.238±0.050 0.145±0.057 0.313±0.054
Safety 0.359±0.118 0.521±0.062 0.442±0.061 0.332±0.069 0.533±0.053
To further investigate the influence of prompt semantics, we randomly rewrite some prompts in the original test dataset, creating new test datasets with prompts different from the downstream dataset. We ask GPT-4o (OpenAI, 2024) to perform two different strategies on these prompts, as detailed in Appendix 8.2. One strategy alters the expression without changing the meaning, and the other generates a new prompt within the same category. As shown in Figure 5, we find that the correlation is less affected by paraphrasing under BoN. However, the correlation measured by the Spearman coefficient on PPO continuously weakens as we paraphrase a larger ratio of prompts.
(a) The correlation trend on BoN.
(b) The correlation trend on PPO.
Figure 5: The correlation between policy regret and accuracy on datasets with different ratios of paraphrased prompts evaluated by the Spearman coefficient and MRR. Different line styles are used to represent datasets paraphrased by different strategies.
Finding 4:
Increasing the number of responses per prompt can enhance the correlation between measured RM error and policy regret.
We first examine the strategy of response expansion per prompt by preparing $5$ responses per prompt and employing various metrics. The details about the computation of these metrics can be found in Appendix 8.8. These metrics include common correlation measures, as well as metrics adapted from Information Retrieval. We also explored other metrics for comparison, e.g., expected calibration error and Bo5. Note that some metrics require absolute scoring of responses, which may be inaccessible in practice. As shown in Table 4, metrics evaluated on the dataset with more responses consistently achieve higher correlations in comparison to the accuracy on the dataset with two responses per prompt. Moreover, the $\xi$ correlation (Chatterjee, 2020) generally demonstrated the best performance. The ECE metric, despite its relative commonness, does not show a significant correlation with regret.
Table 4: The correlation between different metrics and policy regret. The Accuracy-pair stands for the baseline that only samples two responses per prompt and computes the accuracy. Metrics that require an absolute score for each response are suffixed with *.
{tblr}
width = colspec = Q[240]Q[156]Q[131]Q[25]Q[156]Q[131], row1 = c, row2 = c, cell11 = r=2, cell12 = c=20.287, cell15 = c=20.287, cell32 = c, cell33 = c, cell34 = c, cell35 = c, cell36 = c, cell42 = c, cell43 = c, cell44 = c, cell45 = c, cell46 = c, cell52 = c, cell53 = c, cell54 = c, cell55 = c, cell56 = c, cell62 = c, cell63 = c, cell64 = c, cell65 = c, cell66 = c, cell72 = c, cell73 = c, cell74 = c, cell75 = c, cell76 = c, cell82 = c, cell83 = c, cell84 = c, cell85 = c, cell86 = c, cell92 = c, cell93 = c, cell94 = c, cell95 = c, cell96 = c, cell102 = c, cell103 = c, cell104 = c, cell105 = c, cell106 = c, cell112 = c, cell113 = c, cell114 = c, cell115 = c, cell116 = c, cell122 = c, cell123 = c, cell124 = c, cell125 = c, cell126 = c, hline1,13 = -0.08em, hline2 = 2-3,5-6, hline3,12 = -, Metrics & BoN PPO
Spearman corr. MRR Spearman corr. MRR
Pearson corr.* 0.663±0.012 0.657±0.040 0.685±0.016 0.545±0.020
Spearman corr. 0.664±0.015 0.664±0.046 0.680±0.019 0.552±0.030
Kendall $\tau$ corr. 0.665±0.016 0.667±0.042 0.686±0.019 0.558±0.035
Accuracy 0.656±0.015 0.648±0.036 0.671±0.018 0.554±0.025
Bo5* 0.666±0.020 0.663±0.058 0.670±0.024 0.543±0.032
ECE 0.173±0.016 0.364±0.027 0.063±0.014 0.382±0.022
MRR 0.650±0.018 0.646±0.054 0.675±0.022 0.565±0.042
NDCG 0.655±0.032 0.614±0.071 0.658±0.039 0.562±0.053
$\xi$  corr. 0.677±0.026 0.649±0.051 0.688±0.026 0.583±0.047
Accuracy-pair 0.635±0.039 0.586±0.080 0.642±0.047 0.567±0.082
Although measuring more responses per prompt yields a higher correlation, it can be unfair without controlling for the total number of samples (prompt-response pairs) in the test dataset. Thus, we further explore the question of whether we should increase the number of responses or prompts while maintaining a constant sample size. We fix the sample size in the test dataset while varying the number of responses. As shown in Figure 6(a) and Figure 6(b), it is generally more effective to increase the number of responses rather than the number of prompts. However, this advantage decreases with the growing sample size, possibly due to reaching an upper bound in the correlation between regret and measured accuracy.
(a) The correlation trend on BoN.
(b) The correlation trend on PPO.
Figure 6: The Spearman coefficient between accuracy and policy regret with test datasets having the same number of samples but varying numbers of responses per prompt.
The annotation cost can be another important factor since building a test dataset with more than two responses can be tricky. To explore this factor, we further examine the question that should we increase responses or prompts under a constant annotation budget. To conduct the experiment, we consider the noise in the annotation process. Among various probability ranking models (Critchlow et al., 1991), we adopt the most commonly used one, i.e., the Bradley-Terry model (Bradley & Terry, 1952). Given a prompt $x$ and two responses $y_{i}$ and $y_{j}$, the probability that $y_{i}$ is preferred to $y_{j}$ is computed by:
$P(y_{i}\succ y_{j}|x)=\frac{1}{1+e^{(r_{j}-r_{i})/\beta}}$ (5)
where $r_{i}$ and $r_{j}$ is the golden reward score given to the response $y_{i}$ and $y_{j}$ in our case. The parameter $\beta$, which can be seen as an indicator of the annotation quality, is set to $0.5$. We evaluate the annotation cost by measuring the number of pairwise comparisons. This setting fits the Bradley-Terry model and is also commonly adopted in real-world practice. The annotation process is stimulated by performing a sorting algorithm based on pair-wise comparison. The final results are presented in Figures 7(a) and 7(b). We find that under the BoN setting, including more responses is more beneficial with the same annotation budget. However, with a limited annotation budget, the benefits quickly diminish as the number of responses grows. Additionally, in the PPO setting, increasing the number of responses does not yield significant advantages.
(a) The correlation trend on BoN.
(b) The correlation trend on PPO.
Figure 7: The Spearman coefficient between policy regret and accuracy on test datasets with the same annotation budget but varying the number of responses per prompt.
5 What’s the relationship between RM error and Policy Regret?
In this section, we investigate the relationship between RM error and policy regret. The translation from RM error to the policy regret can be seen as the result of the reward model overoptimization. Therefore, to accurately predict policy regret, the quantified RM error should be able to reveal the potential overoptimization. We first theoretically explore this issue. Then we empirically examine the relationship and analyze the observed optimization dynamics.
Finding 5:
Accuracy alone can be insufficient to capture the potential RM overoptimization.
Goodhart’s Law (Goodhart, 1984) is the main cause for the translation from RM error to policy regret. It states that optimizing a less effective metric rather than the golden metric leads to system failure. Such a phenomenon is also commonly termed reward model overoptimization (Gao et al., 2022) under the context of RLHF. Among the various kinds of Goodhart’s effect that lead to reward model overoptimization, the Regressional Goodhart effect (Manheim & Garrabrant, 2019) is the most fundamental and unavoidable. It occurs when the proxy reward model $r$ is essentially the golden reward function $r^{*}$ mixed with some independent noise $z$, namely:
$r=r^{*}+z$ (6)
Assuming that only the Regressional Goodhart effect occurs and both the reward score and noise are normally distributed, i.e. $r^{*}\sim\mathcal{N}(0,\sigma_{r}^{2})$ and $z\sim\mathcal{N}(0,\sigma^{2})$, the relationship between accuracy $d_{r}^{acc}$ and the degree of overoptimization $d_{\pi}$  [^2] The definition of $d_{\pi}$ and detailed derivation are provided in Appendix 8.9. can be expressed by the following parametric equations:
[…]
where $\Phi$ denotes the cumulative distribution function of the standard normal distribution.. The equation is too complex to be solved analytically, so we resort to a numerical solution. We plot the expected relationship between accuracy and the degree of overoptimization in Figure 8(a).
(a) Expected trend of accuracy $d_{r}^{acc}$ and the degree of overoptimization $d_{\pi}$ by varying the noise $\sigma$.
(b) The trend of the accuracy $d_{r}^{acc}$ and the estimated degree of overoptimization $d_{\pi}$ under BoN.
Figure 8: Trends between the accuracy $d_{r}^{acc}$ and the degree of overoptimization $d_{\pi}$.
However, the trend under the BoN setting, as plotted in Figure 8(b), reveals outliers beyond the expected relationship. These discrepancies may stem from other forms of Goodhart’s effects (Manheim & Garrabrant, 2019) beyond the Regressional type. Figure 9 illustrates the optimization dynamics when using different golden RMs and optimizing towards various proxy RMs. We observe distinct overoptimization phenomena for golden-proxy RM pairs that have similar accuracy levels. In Figure 9(a), there is nearly no drop in golden reward scores as the KL divergence increases, indicating that the Regressional Goodhart’s effect is dominant in this case. In contrast, Figure 9(b) shows a noticeable decline in golden reward scores across nearly all proxy RMs, suggesting the influence of more complex Goodhart’s effects. These differences suggest that solely relying on accuracy may be insufficient to predict potential overoptimization.
(a) RM with $\alpha=0$ as golden reward model.
(b) RM with $\alpha=0.15$ as golden reward model.
Figure 9: Overoptimization trends under BoN setting. The change of golden reward score when optimizing towards different proxy RMs are plotted. The distance $d$ is computed as $\sqrt{D_{KL}(\pi||\pi_{0})}$ (Bai et al., 2022), where the KL divergence of best-of-$n$ sampling can be computed by $\log n-\frac{n-1}{n}$ (Hilton, 2023). The rewards are computed by an unbiased estimator (Nakano et al., 2022).
6 Related Works
Reinforcement Learning from Human Feedback (RLHF) (Bai et al., 2022; Ouyang et al., 2022; Stiennon et al., 2022) has been a common strategy for the alignment of Large Language Models (Yang et al., 2024; Dubey et al., 2024), in which the reward model plays a crucial role. However, this methodology faces several challenges (Casper et al., 2023; Lang et al., 2024; Armstrong & Mindermann, 2019; Skalse et al., 2022), such as reward model overoptimization (Gao et al., 2022; Lehman et al., 2019), where optimizing policy towards proxy reward model may lead to the ground truth performance deterioration. This phenomenon, as a special case of reward hacking (Krakovna, 2020), can seen as the result of Goodhart’s law (Goodhart, 1984; Zhuang & Hadfield-Menell, 2021).
The RM as an imperfect proxy of the golden preference necessitates the evaluation of whether the reward model accurately captures the real preference (Michaud et al., 2020; Russell & Santos, 2019). Many works consider the measurements to predict the potential consequence due to the difference of reward functions (Ng et al., 1999; Skalse et al., 2023). (Gleave et al., 2021; Skalse et al., 2024) propose the metric quantify the difference between the reward functions and induce regret bounds for optimal policies. However, some works (Zhuang & Hadfield-Menell, 2021; Fluri et al., 2024) propose negative visions of optimization towards imperfect proxy RMs. These works assume the accessibility of the golden reward function, which can be false under the RLHF setting. The common practice under RLHF is to compute accuracy on a fixed preference dataset (Lambert et al., 2024), which remains the question about theoretical or empirical validation of performance prediction (Zhou et al., 2024; Kim et al., 2024).
7 Conclusion
Our study highlights that although there is a weak positive correlation between accuracy and policy performance, RMs with similar accuracy can result in varying policy outcomes. Moreover, the method used to measure accuracy significantly impacts prediction performance. Finally, we find that accuracy alone may not fully reflect the phenomenon of overoptimization. Overall, we suggest a more cautious attitude about RM performance as indicated by accuracy and emphasize the importance of developing more sophisticated RM evaluation techniques..